Help ercises Ida Sidha Karya Company is a family owned company located in the village of Glanyar on the Island of Ball in Indonesia. The company produces a handcrafted Ballnese musical Instrument called a gamelan that is similar to a xylophone. The gamelans are sold for $870. Selected data for the companys operations last year follow Unita is beginning Ieventory 30 Variable costs per unitr 105 Variable selling and adsinistrative 65,000 s 21,000 ixed selling and adainistrative The absorption costing income statement prepared by the companys accountant for last year appears below 200,100 Cost of goods nold Be11ing and administrative expense Net operating income .900 1 Under absorption costing, how much fixed manufacturing overhead cost is included in the companys inventory at the end of last year? 2. Prepare an income statement for last year using variable costing Complete this question by entering your answers in the tabs below < Prev 2 of 5 Next > 8.
media%2F55e%2F55e9bfbe-acba-49e3-89f5-4d
media%2Fe47%2Fe477fe87-9a23-4b32-b96a-6f
Answer #1
Solution:
1. Fixed manufacturing overhead cost included in inventory $7,500
Working Notes:
Fixed manufacturing overhead per unit = Total fixed manufacturing overhead/total units produced
=$65,000/260 units
=$250 per unit
Units in ending inventory = 30 units
Fixed manufacturing overhead cost included in inventory = units in ending inventory x Fixed manufacturing overhead cost per unit
=30 units x $250 per unit
=$7,500
2.
Ida Sidha Karya Company
Variable Costing Income Statement
Sales $200,100
Variable expenses:
Variable cost of goods sold $109,250
Variable selling and administrative expense $3,450
$112,700
Contribution margin $87,400
Fixed expenses:
Fixed manufacturing overhead $65,000
Fixed selling and administrative $21,000
$86,000
Net operating income $1,400
Working Notes:
Follow the calculation show in alphabets in 1st column to understand how the amounts are calculated.
Ida Sidha Karya Company
Variable Costing Income Statement
a Sales $200,100
[230 units sold x $870 selling price per unit]
Variable expenses:
b Variable cost of goods sold 109,250
[230 units sold x $475 Variable cost of goods sold per units ]
c Variable selling and administrative expense 3,450
[230 units sold x $15 Variable selling and admin. per units ]
d=b+c 112,700
e=a-d Contribution margin 87,400
Fixed expenses:
f Fixed manufacturing overhead 65,000
g Fixed selling and administrative 21,000
h=f+g 86,000
i=e-h Net operating income 1,400
Notes: Variable cost of goods sold per units  
Variable costs per unit:
       Direct materials 105
       Direct labor 325
       Variable manufacturing overhead 45
Variable cost of goods sold per units   475
Please feel free to ask if anything about above solution in comment section of the question.